A car starting from rest passes two successive milestones A and B that are 250 m apart. It takes 60 s to pass A and another 60 s to pass from A to B. Find the velocities of the car when it passes milestone A and milestone B. Assume that the acceleration of the car is uniform 


1
Expert's answer
2022-03-02T14:38:14-0500

"x=\frac{at^2}2,"

"x+s=\frac{a(2t)^2}2=2at^2,"

"s=\frac{3at^2}2,"

"a=\frac{2s}{3t^2}=0.046~\frac m{s^2},"

"v_1=at=2 .8~\frac ms,"

"v_2=2at=5.6~\frac ms ."